OpenStudy (anonymous):

Find the product. (9z + 11)^2 a. 81z^2 + 198z + 121 b. 9z^2 + 198z + 121 c. 81z^2 + 121 d. 9z^2 + 121

OpenStudy (anonymous):

a

jimthompson5910 (jim_thompson5910):

(9z + 11)^2 (9z + 11)(9z + 11) 9z(9z + 11)+11(9z + 11) From here, distribute and simplify

OpenStudy (callisto):

(a + b)^2 = a^2 + 2ab + b^2 Put a =9z and b =11 What do you get?